# Simulating Quantum Circuits with Tree Tensor Networks using Density-Matrix Renormalization Group Algorithm

### v2

@@ Abstract
Quantum computing offers the potential for computational abilities that can go beyond classical machines. However, they are still limited by several challenges such as noise, decoherence, and gate errors. As a result, efficient classical simulation of quantum circuits is vital not only for validating and benchmarking quantum hardware but also for gaining deeper insights into the behavior of quantum algorithms. A promising framework for classical simulation is provided by tensor networks. Recently, the Density-Matrix Renormalization Group (DMRG) algorithm was developed for simulating quantum circuits using matrix product states (MPS). Although MPS is efficient for representing quantum states with one-dimensional correlation structures, the fixed linear geometry restricts the expressive power of the MPS. In this work, we extend the DMRG algorithm for simulating quantum circuits to tree tensor networks (TTNs). The framework allows us to preserve the TTN structure upon application of gates. To benchmark the method, we simulate random and QAOA circuits with various two-qubit gate connectivities. For the random circuits, we devise tree-like gate layouts that are suitable for TTN and show that TTN requires less memory than MPS for the simulations. For the QAOA circuits, a TTN construction that exploits graph structure significantly improves the simulation fidelities. Our findings show that TTNs provide a promising framework for simulating quantum circuits, particularly when gate connectivities exhibit clustering or a hierarchical structure.
@@ I Introduction
Quantum computing promises revolutionary computational capabilities, offering the potential to solve problems that are intractable for classical machines [1, 2, 3]. However, despite the allure of exponentially faster algorithms, quantum computers today are still hampered by several challenges [4]. Current devices are limited by noise, decoherence, and gate errors, all of which constrain their reliability. As a result, efficient classical simulation of quantum systems is vital for validating and benchmarking quantum hardware and gaining deeper insights into the behavior of quantum algorithms [5, 6, 7].
@@
Tree tensor networks have emerged as a powerful tool in both many-body physics and quantum chemistry, with applications spanning ground state calculations [32, 33, 34, 35, 36, 37, 38], real-time dynamics [39, 40, 41, 42, 43, 44, 45], and the study of complex molecular systems [46, 47, 48, 49]. However, TTNs have not been explored extensively in the context of quantum circuit simulation. Their structural flexibility allows for adapting the ansatz to match the underlying connectivity or correlation patterns of the quantum system. A direct application of gates on a TTN leads to a complex tensor network structure, which would then require transformation back into the TTN shape. However, naive transformation of a tensor network can result in an exponentially large memory requirement, scaling with the number of qubits. In this work, we extend the DMRG algorithm for simulating quantum circuits to TTNs. The framework allows us to preserves the shape of the TTN upon the application of gates. It also enables dynamical change of the TTN structure at intermediate stages during the execution of the circuit. The key implication of our work is that TTNs provide a promising framework for simulating quantum circuits, especially when gate connectivities exhibit some form of clustering or hierarchical structure. Therefore, the flexibility of TTNs allows them to adapt to a wide range of quantum circuit architectures. We focus on random circuits and QAOA circuits with various connectivities, including a special class of circuits referred to as tree-like circuits. For tree-like random circuits, TTNs require less memory than MPS for comparable accuracy due to their ability to better capture long-range correlations. In QAOA simulations for MaxCut on 3-regular and tree-like connected graphs, a naive TTN construction exploiting graph structure significantly improves fidelity compared to a blind approach. We achieve near-perfect fidelity for structured bridged 3-regular and tree-like connected graphs despite having small bond dimensions.
@@
The paper is organized as follows. In Sec. II.1, we introduce the TTN ansatz and outline its key features. Section II.2 presents the DMRG algorithm as applied to the simulation of quantum circuits. In Sec. III, we report our results for the simulation of both random and QAOA circuits across various qubit connectivity configurations. Finally, in Sec. IV we provide our conclusions and outlook.
@@ II.1 Tree tensor network
where the summation is over all the virtual indices collectively denoted by $\{\chi\}$. $T_{\{\chi\}_{v}}^{[v]}$ is the tensor associated with the node $v$ on the tree network, $\{\chi\}_{v}$ is the set of virtual indices that are connected to the node and $\{q\}_{v}$ represents the set of physical indices that are connected to the node. Note that $\{q\}_{v}$ can be an empty set, whereas $\{\chi\}_{v}$ is always non-empty as every node in the TTN is connected to the rest of the network through its virtual bonds. In this study, the total number of indices (i.e., elements of $\{q\}_{v}\cup{\{\chi\}}_{v}$) for each tensor in the TTNs is at most five.
TTNs provide a loop-free tensor network representation that allows full exploitation of tensor network gauge freedom. Smart gauging can significantly reduce the number of required tensor contractions and thereby improve computational efficiency [50, 51]. This feature is crucial for the DMRG algorithm used in this study as it greatly simplifies the compression step, discussed in Sec. II.2. In contrast to MPS, the distance between physical indices in a TTN scales logarithmically with the system size as $\mathcal{O}(\log N)$, whereas in MPS, it scales linearly $\mathcal{O}(N)$. Since connected correlation functions typically decay exponentially with the path length in a tensor network, the shorter paths in TTNs allow for a more faithful representation of long-range correlations compared to MPS [29, 32]. This property makes TTNs more efficient at representing systems with long-range interactions or quantum circuits with non-local gates. In this work, we leverage this flexibility to simulate a wide class of circuits with a carefully constructed TTN.
@@ II.2 Quantum circuit simulation
In the DMRG algorithm for simulating quantum circuits, the main idea is to approximate the evolving quantum state using a TTN ansatz as illustrated in Fig. 2(a). The approximation is performed through iterative compression, where one tensor is locally optimized at a time [8]. At no stage is the full evolved quantum state computed explicitly during this process. Starting with a TTN at step $i$ denoted by |$\ket{\Psi^{i}}$, this ideally approximates the exact quantum state after the application of $i$ gates. When a sequence of $k$ additional gates, $\hat{U}_{i+k}\cdots\hat{U}_{i+1}$, is applied, the updated state is given as
@@
where the Tr denotes the full contraction of the network with summation over all indices. The optimal target tensor (up to normalization) is then obtained from the environment tensor $F^{[v]}$, which is computed by contracting the overlap network $\braket{\widetilde{\Psi}^{i+k}}{\Psi^{i+k}}$ without the tensor $\widetilde{T}^{[v]*}$, as shown in Fig 2(c–d). The resulting expression for the optimal target tensor is given as
@@ III.1 Random circuits
Although random circuits are not typically designed to solve practical computational problems, they play a crucial role as benchmarks for demonstrating quantum advantage on near-term quantum devices [55, 56, 52]. Their significance lies in the high degree of entanglement they generate, which poses a major challenge for classical simulation methods such as tensor networks that rely on low-entanglement structures. In these simulations, we focus on the error per gate $\epsilon$, a metric commonly employed for benchmarking NISQ devices [52, 53, 54]. A very recent preprint by Quantinuum demonstrated random circuit implementations on arbitrary geometries [57] which they benchmarked using the DMRG algorithm with MPS using error per gate $\epsilon$ as the metric. However, the linear structure of MPS limits its expressive power for such circuits. In this context, TTN may offer a more suitable framework for benchmarking these devices.
@@ III.2 QAOA
To test our method on a useful task, we simulate the QAOA circuit on various graphs for the MaxCut problem. The MaxCut problem involves partitioning the vertices of a graph into two subsets to maximize the number of edges crossing between these subsets. The QAOA circuit for the MaxCut problem on a graph with $N$ vertices with depth $p$ and $2p$ parameters is expressed as
@@
For random 3-regular graphs, we employ a “naive” clustering approach to construct the TTN. Clusters are first identified within the graph such that there is a high density of edges within each cluster and only a few edges connecting different clusters. This clustering is achieved using the algorithm described in [58], which is designed for community detection in networks. Once the clusters are identified, qubits belonging to the same cluster are positioned closer together in the TTN. However, this approach does not account for inter-cluster relationships, which limits its ability to fully exploit the graph’s connectivity. We compare this naive approach to a “blind” one that doesn’t use any structural information about the graph during TTN construction. Despite its simplicity, the naive method demonstrates significant improvements in fidelity compared to the blind approach, as shown in Fig. 10. The results are presented for the simulation of a QAOA circuit with 64 qubits at depth $p=1$ on random 3-regular graphs. The naive approach achieves better results with $\chi=16$ than the blind approach with $\chi=64$, highlighting the effectiveness of leveraging even basic structural information in TTN construction can enhance algorithm performance. These findings suggest that further improvements in fidelity can be achieved with more sophisticated TTN constructions that exploit both intra- and inter-cluster relationships.
@@ IV Conclusion and Outlook
Tensor networks offer a ideal framework for classical simulation of quantum circuits and have been successfully applied in various contexts [5, 28, 29, 7, 31, 6, 25, 27]. In particular, the DMRG algorithm was generalized to simulate quantum circuits with a finite fidelity using matrix product states [26]. However, despite their efficiency, MPS are fundamentally limited by their chain-like structure. This structural constraint restricts their ability to represent quantum states with more complex entanglement patterns. To address this limitation, we extend the DMRG algorithm for quantum circuit simulation to tree tensor networks.
@@
In this study, we considered the simulation of random circuits and QAOA circuits across various two-qubit gate connectivities, including a special class of circuits referred to as tree-like circuits. The key implication of our work is that TTNs provide a promising framework for simulating quantum circuits, especially when gate connectivities exhibit some form of clustering or hierarchical structure. For the tree-like circuits with random gates, we showed that MPS requires substantially more memory than TTN to achieve comparable error per gate. This advantage is due to TTN’s ability to better capture long-range correlations than MPS. For QAOA circuits on random 3-regular graphs, we demonstrated that a naive TTN construction approach based on clustering graph nodes greatly improves fidelity compared to a blind approach that ignores structural information. To further emphasize the importance of exploiting structure, we simulated the QAOA circuit for bridged 3-regular graphs and graphs with two-qubit gate connectivity that aligns with the structure of tree-like circuits. For both types of graphs, we achieved high fidelities at very small bond dimensions. Our results also highlight that circuits designed for practical tasks, such as QAOA, are significantly more amenable to compression than random circuits.
@@
Our findings demonstrate that tree tensor networks offer an efficient and flexible framework for simulating quantum circuits. A natural extension of our work is to consider a two-site version of the compression step. Although the previous work using MPS reported no significant improvement from the two-site variant over the single-site version [26], its potential advantages in the TTN setting remain worth investigating. Further performance gains could be achieved by constructing optimal TTNs using graph embedding techniques or enabling adaptive changes to the TTN structure during simulation, as developed for MPS in [59]. Notably, the compression step in our algorithm is quite general and can be incorporated into other TTN-based quantum circuit algorithms, such as those proposed in Ref. [29]. Recent works, such as Quantinuum’s realization of random quantum circuits in arbitrary geometries and benchmarking them with MPS [57], highlight the relevance of our approach as an alternative benchmarking tool. The framework employed here can also be extended to simulate the time evolution of a quantum many-body system [60]. Finally, for certain applications like QAOA circuits, where the entanglement generated remains relatively low [61], our method offers a particularly efficient and practical solution.

### v3

@@ Abstract
Quantum computing offers the potential for computational abilities that can go beyond classical machines. However, they are still limited by several challenges such as noise, decoherence, and gate errors. As a result, efficient classical simulation of quantum circuits is vital not only for validating and benchmarking quantum hardware but also for gaining deeper insights into the behavior of quantum algorithms. A promising framework for classical simulation is provided by tensor networks. Recently, the Density-Matrix Renormalization Group (DMRG) algorithm was developed for simulating quantum circuits using matrix product states (MPS). Although MPS is efficient for representing quantum states with one-dimensional correlation structures, the fixed linear geometry restricts the expressive power of the MPS. In this work, we extend the DMRG algorithm for simulating quantum circuits to tree tensor networks (TTNs). The framework employs a variational compression scheme that optimizes the TTN to approximate the evolved quantum state. To benchmark the method, we simulate random circuits and the quantum approximate optimization algorithm (QAOA) with various two-qubit gate connectivities. For the random circuits, we devise tree-like gate layouts that are suitable for TTN and show that TTN requires less memory than MPS for the simulations. For the QAOA circuits, a naive TTN construction that exploits graph structure significantly improves the simulation fidelities. Our findings show that the DMRG algorithm with TTNs provides a promising framework for simulating quantum circuits, particularly when gate connectivities exhibit clustering or a hierarchical structure.
@@ I Introduction
Quantum computing promises revolutionary computational capabilities, offering the potential to solve problems that are intractable for classical machines [1, 2, 3]. However, despite the allure of exponentially faster algorithms, quantum computers today are still hampered by several challenges [4, 5]. Current devices are limited by noise, decoherence, and gate errors, all of which constrain their reliability. As a result, efficient classical simulation of quantum systems is vital for validating and benchmarking quantum hardware and gaining deeper insights into the behavior of quantum algorithms [6, 7, 8].
@@
Tree tensor networks have emerged as a powerful tool in both many-body physics and quantum chemistry, with applications spanning ground state calculations [33, 34, 35, 36, 37, 38, 39], real-time dynamics [40, 41, 42, 43, 44, 45, 46, 47], and the study of complex molecular systems [48, 49, 50, 51]. However, TTNs have not been explored extensively in the context of quantum circuit simulation. Their structural flexibility allows for adapting the ansatz to match the underlying connectivity or correlation patterns of the quantum system. A direct application of gates on a TTN leads to a complex tensor network structure, which would then require transformation back into the TTN shape. However, naive transformation of a tensor network can result in an exponentially large memory requirement, scaling with the number of qubits. Ref. [30] addresses this challenge by introducing the concept of threading entanglement through the tree. This method is equivalent to approximating the true output quantum state using a truncation method based on singular value decomposition (SVD). The truncation is based on discarding small singular values which can introduce a notable error with each gate. These errors compound with subsequent gate applications, potentially leading to a significant loss of fidelity.
In this work, we extend the Density-Matrix Renormalization Group (DMRG) algorithm, originally developed for simulating quantum circuits with MPS [27], to TTNs. The central idea of this method is a variational compression scheme that optimizes the TTN to approximate the evolved quantum state. It iteratively optimizes each tensor to find the best possible TTN representation of the evolved quantum state for a given bond dimension. This variational approach is superior to the SVD-based compression method as it minimizes the distance to the target state and avoids the compounding errors associated with successive SVD truncations [9, 27]. Furthermore, the framework preserves the shape of the TTN upon the application of gates and enables dynamical change of the TTN structure mid-simulation. The key implication of our work is that the DMRG algorithm with TTNs provides a promising framework for simulating quantum circuits, especially when gate connectivities exhibit some form of clustering or hierarchical structure. Therefore, the flexibility of TTNs allows them to adapt to a wide range of quantum circuit architectures. We focus on random circuits and QAOA circuits with various connectivities, including a special class of circuits referred to as tree-like circuits. For tree-like random circuits, TTNs require less memory than MPS for comparable accuracy due to their ability to better capture long-range correlations. In QAOA simulations for MaxCut on 3-regular and tree-like connected graphs, a naive TTN construction exploiting graph structure significantly improves fidelity compared to a blind approach. We achieve near-perfect fidelity for structured bridged 3-regular and tree-like connected graphs despite having small bond dimensions.
@@
The paper is organized as follows. In Sec. II.1, we introduce the TTN ansatz and outline its key features. Section II.2 presents the DMRG algorithm as applied to the simulation of quantum circuits. In Sec. III, we report our results for the simulation of both random and QAOA circuits across various qubit connectivity configurations. Finally, in Sec. IV we provide our conclusions and outlook. Appendix A presents a comparison between the DMRG algorithm and the SVD-based truncation scheme for simulating quantum circuits, and Appendix B provides a comparison between MPS and TTN within the DMRG framework.
@@ II.1 Tree tensor network
where the summation is over all the virtual indices collectively denoted by $\{\chi\}$. $T_{\{q\}_{v},\{\chi\}_{v}}^{[v]}$ is the tensor associated with the node $v$ on the tree network, $\{\chi\}_{v}$ is the set of virtual indices that are connected to the node and $\{q\}_{v}$ represents the set of physical indices that are connected to the node. Note that $\{q\}_{v}$ can be an empty set, whereas $\{\chi\}_{v}$ is always non-empty as every node in the TTN is connected to the rest of the network through its virtual bonds. In this study, we employ layer-wise regular tree structures for the TTN ansatz. Specifically, all non-leaf nodes at a given layer have the same number of children, though this number can vary across layers. This provides more flexibility than a strictly regular tree, where every non-leaf node must have the same number of children. The total number of indices (i.e., elements of $\{q\}_{v}\cup{\{\chi\}}_{v}$) for each tensor in the TTNs is at most five with at most four virtual indices. Conventionally, TTNs are built using tensors with at most three virtual legs to control scaling with the bond dimension [52]. We relax this restriction for two main reasons. First, such a constraint severely limits the flexibility in choosing tree topologies for arbitrary qubit counts. For instance, creating a layer-wise regular TTN for a 27-qubit system becomes impractical under this condition. Second, when the qubit connectivity exhibits a hierarchical clustering, it is beneficial to capture that structure at higher levels of the tree. This may require tensors with more than three virtual legs to efficiently group strongly correlated qubits. Such clustering helps keep bond dimensions manageable, which in turn can mitigate the additional scaling overhead introduced by allowing more virtual legs.
TTNs provide a loop-free tensor network representation that allows full exploitation of tensor network gauge freedom. This gauge freedom allows the TTN to be brought into the so-called canonical form [53, 52]. In this form, one tensor is designated as the orthogonality center, and all other tensors are isometric with respect to the direction pointing toward it. Smart gauging can significantly reduce the number of required tensor contractions and thereby improve computational efficiency. For example, the norm of the TTN reduces to a single contraction of the orthogonality center with its conjugate. This feature is crucial for the DMRG algorithm used in this study as it greatly simplifies the compression step, as discussed in Sec. II.2. In contrast to MPS, the distance between physical indices in a TTN scales logarithmically with the system size as $\mathcal{O}(\log N)$, whereas in MPS, it scales linearly $\mathcal{O}(N)$. Since connected correlation functions typically decay exponentially with the path length in a tensor network, the shorter paths in TTNs allow for a more faithful representation of long-range correlations compared to MPS [30, 33]. This property makes TTNs more efficient at representing systems with long-range interactions or quantum circuits with non-local gates. In this work, we leverage this flexibility to simulate a wide class of circuits with a carefully constructed TTN.
@@ II.2 Quantum circuit simulation
In the DMRG algorithm for simulating quantum circuits, the main idea is to approximate the evolving quantum state using a TTN ansatz as illustrated in Fig. 2(a). The approximation is performed through a variational compression scheme by iteratively optimizing individual tensors to minimize the distance to the target state. This approach is fundamentally superior to the SVD-based method as it avoids the accumulation of truncation errors [9, 27]. We demonstrate this explicitly in Appendix A by comparing our approach with the one given in [30].
To formalize our algorithm, let us consider a TTN at step $i$ denoted by |$\ket{\Psi^{i}}$, which ideally approximates the exact quantum state after the application of $i$ gates. When a sequence of $k$ additional gates, $\hat{U}_{i+k}\cdots\hat{U}_{i+1}$, is applied, the updated state is given as
@@
where the Tr denotes the full contraction of the network with summation over all indices. The optimal target tensor (up to normalization) is then obtained from the environment tensor $F^{[v]}$, which is computed by contracting the overlap network $\braket{\widetilde{\Psi}^{i+k}}{\Psi^{i+k}}$ without the tensor $\widetilde{T}^{[v]*}$ [9], as shown in Fig 2(c–d). The resulting expression for the optimal target tensor is given as
@@ III.1 Random circuits
Although random circuits are not typically designed to solve practical computational problems, they play a crucial role as benchmarks for demonstrating quantum advantage on near-term quantum devices [57, 58, 54]. Their significance lies in the high degree of entanglement they generate, which poses a major challenge for classical simulation methods such as tensor networks that rely on low-entanglement structures. In these simulations, we focus on the error per gate $\epsilon$ as defined in Eq. 9. A very recent work by Quantinuum demonstrated random circuit implementations on arbitrary geometries [59], which they benchmarked using the DMRG algorithm with MPS using error per gate $\epsilon$ as the metric. However, the linear structure of MPS limits its expressive power for such circuits. In this context, TTN may offer a more suitable framework for benchmarking these devices.
@@ III.2 QAOA
To test our method on a useful task, we simulate the quantum approximate optimization algorithm (QAOA) [60, 61, 62, 63]. QAOA is a hybrid quantum-classical variational algorithm designed to tackle combinatorial optimization problems. In the QAOA framework, the optimization task is mapped to a physical system by encoding the problem’s cost function into a Hamiltonian whose ground state represents the optimal solution. The algorithm proceeds by preparing a parameterized quantum state and using a classical optimizer to adjust the parameters in order to minimize the energy of the Hamiltonian.
In this section, we focus on the QAOA circuit on various graphs for the MaxCut problem. The MaxCut problem involves partitioning the vertices of a graph into two subsets to maximize the number of edges crossing between these subsets. The QAOA circuit for the MaxCut problem can on a graph with $N$ vertices with depth $p$ and $2p$ parameters [60, 64, 65, 62] is expressed as
@@
For random 3-regular graphs, we employ a “naive” clustering approach to construct the TTN. Clusters are first identified within the graph such that there is a high density of edges within each cluster and only a few edges connecting different clusters. This clustering is achieved using the algorithm described in [66], which is designed for community detection in networks. Once the clusters are identified, qubits belonging to the same cluster are positioned closer together in the TTN. This procedure effectively acts as a qubit reordering strategy which reduces the number of virtual bonds connecting qubits that belong to the same cluster. However, this approach does not account for inter-cluster relationships, which limits its ability to fully exploit the graph’s connectivity. We compare this naive approach to a “blind” one that doesn’t use any structural information about the graph during TTN construction, while maintaining the same TTN architecture. Despite its simplicity, the naive method demonstrates significant improvements in fidelity compared to the blind approach, as shown in Fig. 10. The results are presented for the simulation of a QAOA circuit with 64 qubits at depth $p=1$ on random 3-regular graphs. The naive approach achieves better results with $\chi=16$ than the blind approach with $\chi=64$, highlighting the effectiveness of leveraging even basic structural information in TTN construction can enhance algorithm performance. These findings suggest that further improvements in fidelity can be achieved with more sophisticated TTN constructions that exploit both intra- and inter-cluster relationships.
@@ IV Conclusion and Outlook
Tensor networks offer an ideal framework for classical simulation of quantum circuits and have been successfully applied in various contexts [6, 29, 30, 8, 32, 7, 26, 28]. In particular, the DMRG algorithm was generalized to simulate quantum circuits with a finite fidelity using matrix product states [27]. However, despite their efficiency, MPS are fundamentally limited by their chain-like structure. This structural constraint restricts their ability to represent quantum states with more complex entanglement patterns. To address this limitation, we extend the DMRG algorithm for quantum circuit simulation to tree tensor networks.
@@
In this study, we considered the simulation of random circuits and QAOA circuits across various two-qubit gate connectivities, including a special class of circuits referred to as tree-like circuits. The key implication of our work is that the DMRG aglorithm with TTNs provides a promising framework for simulating quantum circuits, especially when gate connectivities exhibit some form of clustering or hierarchical structure. For the tree-like circuits with random gates, we showed that MPS requires substantially more memory than TTN to achieve comparable error per gate. This advantage is due to TTN’s ability to better capture long-range correlations than MPS. For QAOA circuits on random 3-regular graphs, we demonstrated that a naive TTN construction approach based on clustering graph nodes greatly improves fidelity compared to a blind approach that ignores structural information. To further emphasize the importance of exploiting structure, we simulated the QAOA circuit for bridged 3-regular graphs and graphs with two-qubit gate connectivity that aligns with the structure of tree-like circuits. For both types of graphs, we achieved high fidelities at very small bond dimensions. Our results also highlight that circuits designed for practical tasks, such as QAOA, are significantly more amenable to compression than random circuits.
@@
Our findings demonstrate that the DMRG algorithm with tree tensor networks offers an efficient and flexible framework for simulating quantum circuits. A natural extension of our work is to consider a two-site version of the compression step. Although the previous work using MPS reported no significant improvement from the two-site variant over the single-site version [27], its potential advantages in the TTN setting remain worth investigating. In this work, we employed layer-wise regular tree structures. Future research could explore and compare more general tree topologies by leveraging more advanced TTN constructions [30, 67, 68, 69]. Further performance gains could be achieved by enabling adaptive changes to the TTN structure during simulation, as developed for MPS in [70]. Notably, the compression step in our algorithm is quite general and can be incorporated into other TTN-based quantum circuit algorithms, such as those proposed in Ref. [30]. Recent works, such as Quantinuum’s realization of random quantum circuits in arbitrary geometries and benchmarking them with MPS [59], highlight the relevance of our approach as an alternative benchmarking tool. The framework employed here can also be extended to simulate the time evolution of a quantum many-body system [71]. Finally, for certain applications like QAOA circuits, where the entanglement generated remains relatively low [72], our method offers a particularly efficient and practical solution.
@@ Acknowledgements.
Appendix A Comparison with SVD-based approach
In this section, we compare the DMRG method used in our work with the SVD-based truncation scheme described in Ref. [30]. The SVD-based approach approximates the quantum state by applying an SVD after each gate operation and discarding singular values below a certain threshold. These truncations introduce an error with each gate. These errors compound with subsequent gate applications, potentially leading to a significant loss of fidelity.
In contrast, the DMRG method employs a variational compression scheme. It iteratively optimizes each tensor to find the best possible TTN representation of the evolved quantum state for a given bond dimension. This variational approach is superior to the SVD-based compression method as it minimizes the distance to the target state and avoids the compounding errors associated with successive SVD truncations [9, 27].
Figure 14: Comparison of fidelities $\mathcal{F}$ as a function of bond dimension $\chi$ obtained from the simulation of QAOA circuits on random 3-regular graphs by the DMRG method and SVD-TTN from Ref. [30]. A binary tree structure is employed for representing the quantum state in both cases. No information regarding the graph is exploited. Results are shown for $N=16$ at depth $p=1$ with $N_{2g}=24$ two-qubit gates.
Figure 15: Comparison of fidelities $\mathcal{F}$ as a function of bond dimension $\chi$ obtained from the simulation of QAOA circuits on random 3-regular graphs by the DMRG method and SVD-TTN from Ref. [30]. A binary tree structure is used for both methods. The naive method described in Sec. III.2 is applied for qubit reordering to improve ansatz expressivity. Results are shown for $N=16$ at depth $p=1$ with $N_{2g}=24$ two-qubit gates.
To illustrate these differences quantitatively, we simulate quantum approximate optimization algorithm (QAOA) circuits for the Max-Cut problem on random 3-regular graphs with $N=16$ qubits using both methods. The fidelity $\mathcal{F}$ of each method is computed relative to exact statevector simulations performed using Qiskit [76]. The choice of $N=16$ enables this exact benchmarking. To ensure a fair comparison, both methods are implemented using the same binary TTN, and fidelities are reported as a function of the bond dimension. Results for Ref. [30] were obtained using the publicly available code provided by the authors.
Figure 16: Comparison of error rates $\epsilon$ as a function of bond dimension $\chi$ obtained from the simulation of random nearest-neighbor circuits using MPS and TTN. The results are for a system of $N=32$ qubits with $N_{2g}=310$ two-qubit gates. The inset shows error rates $\epsilon$ as a memory footprint $M$ obtained from the same simulation.
Fig. 14 presents a “blind” comparison, where both methods use a binary TTN ansatz without exploiting any information about the graph structure. Even at a very low bond dimension of $\chi=4$, our DMRG-based method outperforms the SVD-based approach at $\chi=64$, demonstrating the superior accuracy of variational optimization.
[…]
Fig. 15 shows the same comparison, now incorporating a “naive” qubit reordering strategy as Sec. III.2. In both cases, a binary TTN is used, but the strongly correlated qubits are placed together in the TTN. While both approaches benefit from reordering, our DMRG-based algorithm at $\chi=16$ still outperforms Ref. [30] at $\chi=64$. Notably, the performance improvement is more pronounced for Ref. [30]. This improvement indicates that the SVD-based method is more sensitive to TTN structure and less robust in capturing the effect of long-range gates. This further highlights the advantage of our variational framework.
Appendix B Comparison between MPS and TTN
We compare MPS and TTN as variational ansatz within the DMRG algorithm for simulating the quantum circuits. All simulations are performed on a system of $N=32$ qubits. For TTN, we adopt a binary tree structure. It is important to note that a direct comparison between TTN and MPS is inherently challenging. As a result, our analysis primarily focuses on the bond dimension $\chi$, which determines the maximum half-chain entanglement entropy that each ansatz can represent [9]. In some instances, we also provide a comparison of the memory footprint.
For random circuits, we consider nearest-neighbor connectivity among two-qubit gates. The linear connectivity of these circuits aligns naturally with the structure of MPS which makes them ideal for such simulations. In contrast, TTNs impose a hierarchical grouping of qubits that introduces a structural bias that is not compatible with the geometry of these circuits. This mismatch is evident in Fig. 16, where MPS outperforms TTNs. The performance advantage observed for MPS in linear circuits is similar to the advantage TTNs exhibit in tree-like circuit topologies (see Fig. III.1). These results highlight the importance of choosing an ansatz compatible with the circuit structure for efficient simulation.
For QAOA circuits, we consider both tree-like circuits and random 3-regular graphs. The hierarchical structure of tree-like circuits makes them naturally compatible with the TTN ansatz. As shown in Fig. 17, TTN significantly outperforms MPS for these circuits. This outcome is consistent with expectations, as the linear geometry of MPS limits its ability to efficiently capture the entanglement present in hierarchically structured circuits.
[…]
For QAOA circuits on random 3-regular graphs, we apply a naive qubit reordering strategy, as described in Sec. III.2, to both MPS and TTN simulations. As shown in Fig. 18, TTN slightly outperforms MPS in terms of bond dimensions. However, the inset reveals that the overall memory footprints of the two ansatzes are comparable. This similarity may result from the limitations of the naive reordering scheme, which does not fully exploit the structure of the graph. More advanced TTN construction methods could better capture the graph topology and provide a more efficient representation. In contrast, for MPS, performance improvements are limited unless the graph exhibits linear clustering. A more comprehensive study is warranted to fully assess the trade-offs between the two approaches in general graph-based QAOA instances.